OpenStudy (anonymous):

Trigonometry Help!! Use identities to simplify and show work (remember you are NOT solving for x, just simplifying the expression) 1-2sin^2 x/2

OpenStudy (anonymous):

\[\cos x = 1- 2\sin^2 {x \over 2}\]

OpenStudy (anonymous):

Can you please show me the steps you took to find that answer. :)

OpenStudy (anonymous):

shivam are you there? it says your are typing?

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

\[\cos 2x = \cos (x+x) = \cos x \cos x - \sin x \sin x = \cos^2 x - \sin^2 x\] Now write cos^2 x = 1-sin^2 x and you will get your answer

OpenStudy (anonymous):

In your case , It should be cos x = cos (x/2 + x/2)=.......

OpenStudy (anonymous):

oh? what happened to your long explanation?

OpenStudy (anonymous):

That was invalid :D Made some mistake. This is shoreter and easier :D

OpenStudy (anonymous):

*shorter

OpenStudy (anonymous):

hah..i understand this but are you saying the answer it cos^2x or cos x

OpenStudy (anonymous):

@Emilyori , I gave the derivation for cos (2x). Now you try to derive it for cos (x) similarly

OpenStudy (anonymous):

i am stuck so would it be \[\cos x \div2 \cos x \div2 - \sin x sinx \] or does the sin change

OpenStudy (anonymous):

Shivam...I have to go,, but will return soon...please send me a message to my question if you can...thank you

OpenStudy (anonymous):

Make sure you try again by seeing the above example